# Dim and Small Target Detection for Drone Broadcast Frames Based on Time-Frequency Analysis

### v1

@@ IV-B Simulation Results
TABLE III: Evaluation Metrics of Different Algorithms
Algorithm Average IoU Average Precision Average Recall Detection Latency
@@
Figure 17: Detection accuracy and speed of the proposed algorithm versus sampling durations.
V Conclusion
This paper proposed a dim and small target detection algorithm for drone broadcast frames to achieve effective drone RID. The prior knowledge of modulation parameters and frame structures was extracted through time-frequency analysis of the communication protocol. Transmission frequency and signal bandwidth were leveraged to correct the frequency domain parameters of bounding boxes, while ZC sequences and frame length were utilized to correct the time domain detection parameters of broadcast frames termed as dim and small targets. Besides, a segmented energy refinement method was applied to mitigate time domain parameter deviation caused by interference signals with high energy strength. Since the sampling duration determined the occupied regions of broadcast frame and whether it corresponded to a small target, the trade-off between detection accuracy and speed of the proposed algorithm was established to meet different detection requirements. Simulation results demonstrated that the proposed algorithm outperformed existing algorithms in terms of both detection accuracy and speed. And the proposed algorithms also exhibited the robustness across varying flight distances, different flight visual environment, and diverse types of environment noise.

### v2

@@ V-B Simulation Results
TABLE III: Average Evaluation Metrics of Different Algorithms
Algorithm IoU Precision Recall Detection Latency
@@
Figure 17: Detection accuracy and speed versus sampling durations.
The partial payload information of two broadcast frames from different datasets is shown in Fig. 18, and some privacy information is masked. It can be observed that by decoding the broadcast frames, information such as the speed, position, and serial number of drone and pilot can be obtained. Among them, the serial number and the universally unique identifier (UUID) are both uniquely associated with an individual, which provides a key technical solution for drone RID. In addition, 37 broadcast frames from a total of 14 drones of 10 categories in Table I are decoded. Only 1 frame failed in CRC, i.e., the decoding failed, which means the identification, association, and positioning of drone and pilot is achieved with an accuracy of 97.30% based on the proposed algorithm in this paper.
Figure 18: Payload information in two broadcast frames from different datasets.
VI Conclusion
This paper proposed a dim and small target detection algorithm for broadcast frames to achieve effective drone RID. The prior knowledge of modulation parameters and frame structures was established. Transmission frequency and signal bandwidth were leveraged to correct the frequency domain parameters of bounding boxes, while ZC sequences and frame length were utilized to correct the time domain detection parameters of broadcast frames. Besides, a segmented energy refinement method was applied to mitigate time domain parameter deviation caused by interference signals with high energy strength. Since the sampling duration determined the occupied regions of broadcast frame and whether it corresponded to a small target, the trade-off between detection accuracy and speed of the proposed algorithm was established to meet different detection requirements. For the detected broadcast frames, a decoding algorithm was proposed to recover the essential payload information for RID. Simulation results demonstrated that the proposed algorithms outperformed existing algorithms in terms of both detection accuracy and speed. And the proposed algorithm also maintained robust performance across varying flight distances, different flight visual environment, and diverse types of environment noise. The decoding results also indicated that the effective of utilizing broadcast frames to achieve RID.